Sheffield Corporation purchased a new plant asset on April 1, 2017, at a cost of $880,000. It was estimated to have a useful life of 20 years and a residual value of $370,000, a physical life of 30 years, and a salvage value of $0. Sheffield’s accounting period is the calendar year. Sheffield prepares financial statements in accordance with IFRS.

Calculate the depreciation for this asset for 2017 and 2018 using the straight-line method. (Round answers to 0 decimal places, e.g. 5,275.)

Depreciation
2017 $

2018 $

Calculate the depreciation for this asset for 2017 and 2018 using the double-declining-balance method. (Round answers to 0 decimal places, e.g. 5,275.)

Depreciation
2017 $

2018 $

Calculate the depreciation for this asset for 2017 and 2018 using the straight-line method and assuming Sheffield prepares financial statements in accordance with ASPE. (Do not round intermediate calculations and round answers to 0 decimal places, e.g. 5,275.)

Depreciation
2017 $

2018 $

Assume that additional information has been provided relating to the cost ($880,000). There are three components of the plant asset. Components 1, 2, and 3 have costs of $365,000, $234,000, and $281,000, respectively. The useful lives of components 1, 2, and 3 are 25, 20, and 30 years, respectively.

Determine straight-line depreciation expense for 2017 and 2018 for each component under IFRS if the residual value is $100,000 for component 1, $102,000 for component 2, and $168,000 for component 3. (Do not round intermediate calculations and round answers to 0 decimal places, e.g. 5,275.)
Depreciation expense
2017 2018
Component 1 $

$

Component 2 $

$

Component 3 $

$
Answer #1

1) Cost of asset = $880000, Residual Value = $370000, Useful Life = 20 years.

Amount of depreciation to be charged every year = (Cost-Residual value)/Useful life

= (880000-370000)/20

= $25,500

2)

2) Double Declining Balance
Straight Line Depreciation percent 100% in 20 years
or, 5 % per year
Depreciation rate (%) 2*5% = 10 % per year
Depreciation for the year 2017 = 10% of B.V at beginning of the year
= 10% of 880000
= $ 88000
Depreciation for the year 2018 = 10%*(880000-Accumulated depreciation)
= 10%*(888000-88000)
= $ 79200

3)For the straight-line method, under ASPE, depreciation expense is the higher of two amounts: (i) cost less salvage value over the life of the asset, and (ii) cost less residual value over the asset’s useful life.

(i)Cost = 880000, Salvage Value = Nil, Physical Life = 30 years,

Depreciation = (880,000-0)/30

= $ 29,333

(ii)Cost = 880000, Residual  Value = 370000, Physical Life = 20 years,

Depreciation = (880,000-370000)20

= $ 25500

Since, depreciation calculated under method (i) is higher, it'd be the amount of depreciation for the year 2017 and 2018.

4)Under straight line method, depreciation for the 3 components for 2 years would be:

Component 1 Cost $ 365000
Useful Life 25 years
Residual value $ 100000
Depreciation = (365000-100000)/25
= $ 10600
For 2017 = $ 10600
For 2018 = $ 10600
Component 2 Cost $ 234000
Useful Life 20 years
Residual value $ 102000
Depreciation = (234000-102000)/20
= $ 6600
For 2017 = $ 6600
For 2018 = $ 6600
Component 3 Cost $ 281000
Useful Life 30 years
Residual value $ 168000
Depreciation = (281000-168000)/30
= $ 3766.667
For 2017 = $ 3767
For 2018 = $ 3767